# Nonreciprocal heat transport in the Kitaev chiral spin liquid

### v1

@@ Abstract
Nonreciprocal transport, characterized by its direction-selective nature, holds significant potential for applications in various devices. In this study, we investigate nonreciprocal heat transport in Majorana systems, specifically focusing on the Kitaev chiral spin liquid under external magnetic fields. Our theoretical examination focuses on effects of open boundaries in which the Majorana edge modes exist, and the inversion symmetry is broken, which leads to the Dzyaloshinskii-Moriya interaction (DMI). Through perturbation theory, we demonstrate that DMI induces asymmetric hopping, resulting in the asymmetry of the Majorana band. The results of nonreciprocal heat currents are presented for various directions of external magnetic fields, and we discuss the relation between the current and the field-directions. The potential exists to manipulate both of the directions and magnitude of the non-dissipative current by varying external magnetic fields and apply to non-dissipative heat transfer devices.
@@ I Introduction
where the conductivity $\chi$ and the external drive field $\mathcal{D}$ satisfy $\chi_{\xi}=\chi_{-\xi}$ and $\mathcal{D}_{\xi}=-\mathcal{D}_{-\xi}$. In particular, if the even-order conductivity is finite, the current magnitudes $|\mathcal{J}|$ differ between positive and negative directions. Nonreciprocal transport phenomena result from higher order effects, which break Onsager reciprocal relations. Its essence is the difference in group velocities at $\bm{k}$ and $\bm{-k}$ because of inversion symmetry breaking. Nonreciprocal transport phenomena have been extensively studied in diverse materials, such as superconductors [1, 2, 3, 4], antiferromagnets [5, 6], and topological insulators [7, 8]. However, its exploration in Majorana systems remains relatively limited [9].
@@
Several transport phenomena have been explored in the Kitaev chiral spin liquid. The itinerant Majorana fermions in the bulk act as carriers of spin [18, 19, 20, 21] and heat [22, 23]. Recently, the nonreciprocal heat transport in the Kitaev spin liquid by applying the staggard magnetic fields is discussed [9]. Additionally, it has been suggested that the itinerant Majorana fermions at the edge play an important role in the spin current [24]. However, the nonreciprocal heat transport for the case with the open edge is still unclear, where the non-dissipative heat careers exist.
@@
In this paper, we investigate the nonreciprocal heat current, which arises from chiral Majorana edge states, and can be manipulated by varying external magnetic fields. We conduct the calculation for the system which has the thermal gradient with consideration of the open edge [see Fig. 1(a)]. The unit cell is extended for the upper and lower sides to be the open boundaries, shown in Fig. 1(b). We consider Dzyaloshinskii-Moriya interaction (DMI) at edges due to the broken inversion symmetry at open boundaries. Thereby, asymmetric hoppings and asymmetric Majorana bands appear. The heat current is evaluated by using Boltzmann transport theory. As a result, we find that the nonreciprocal heat current depends on the direction of the magnetic fields.
@@ II.1 Kitaev spin liquid with edge
where $S^{\gamma}_{j}$ is the $\gamma~{}(=x,y,z)$ component of an $S=1/2$ spin operator at site $j$. To avoid confusion, we adopt $x$, $y$ and $z$ as the orientations of the spin space, while $X$ and $Y$ denote the orientations of the real space, as shown in Fig. 1. The initial term in Eq. (2) represents the Kitaev interaction, characterized as an Ising-type interaction only for the nearest-neighbor (NN) sites. The notation $\langle jk\rangle_{\gamma}$ indicates the summation over NN sites connected by $\gamma$-bonds. Here, we set the interaction strength $J$ as the unit of energy with $J>0$, but our results are independent of the sign of $J$. The second term in Eq. (2) is a perturbation term induced by the Zeeman coupling $H_{\mathrm{h}}=-\sum_{j}h_{\gamma}S^{\gamma}_{j}$. The mean flux excitation energy is $\Delta$, and we set $\Delta=0.065J$. The coefficient $3!$ is due to the permutation of three spin operators arising from the perturbative processes. The notation $\langle\!\langle jkl\rangle\!\rangle$ denotes the summation exclusively over triplets of sites $j,k,$ and $l$ that are aligned and interconnected. Eq. (2) can be exactly solved using the well-known Majorana representation, $S^{\gamma}_{j}=\frac{i}{2}b^{\gamma}_{j}c_{j}$. Here, both $b_{j}^{\gamma}$ and $c_{j}$ satisfy the Majorana condition $\eta^{\dagger}=\eta$ and $\{\eta_{j}^{\alpha},\eta_{k}^{\beta}\}=\delta_{jk}\delta^{\alpha\beta}$ (with $\eta_{j}^{\alpha}=b_{j}^{x},b_{j}^{y},b_{j}^{z},c$). We start by considering a thermal current in the ground state, i.e., $u_{j,k}=ib^{\gamma}_{j}b^{\gamma}_{k}=1$ ($j\in{\textrm{even sublattice}}$, $k\in{\textrm{odd sublattice}}$) for all bonds. Each of the first and second terms in Eq. (2) transforms into a NN hopping and a next nearest-neighbor (NNN) hopping, respectively. When $h_{x}h_{y}h_{z}\neq 0$, the Chern number takes values $\nu=\pm 1$, reflecting the sing of $h_{x}h_{y}h_{z}$, and the non-dissipative chiral current is realized at the edge.
@@
Moreover, the DMI at the edges is taken into account due to the broken inversion symmetry at these boundaries. DMI is expressed as,
@@
where the summation is restricted to NN sites ($j\in{\textrm{even sublattice}}$, $k\in{\textrm{odd sublattice}}$), with at least one site belonging to the edge. For simplicity, we assume that the inversion symmetry in the bulk is not broken, considering DMI only at the edges. In the second-order perturbation combining DMI and magnetic fields, the effective Majorana Hamiltonian is given as follows:
@@ II.3 Boltzmann transport equation
$v_{k_{X}n}$ represents the group velocity in the $X$ direction ($v_{k_{X}n}=\partial E_{k_{X}n}/{\partial k_{X}}$) and $n$ is the index of Majorana bands. The contributions from both the Majorana edge bands and the bulk are summed up in the second summation of Eq. (12). $\Omega$ denotes the system size, $\Omega=Nk_{\mathrm{mesh}}$, where $k_{\mathrm{mesh}}=1/N_{\mathrm{unit}}$. (A unit of length is a lattice constant.) We adopt the Boltzmann constant value $k_{\mathrm{B}}=1$, and $f$ represents Fermi distribution function, $f(E_{k_{X}n})=(\exp(E_{k_{X}n}/T)+1)^{-1}$. Here, $\tau$ is the relaxation time. Assuming that backscattering of Majorana edge modes occurs mediated by the bulk excitations at finite temperature, we employ the relaxation time approximation and set $J\tau=1$. The second-order conductivity is nonlinear conductivity, which provides the nonreciprocal contribution.
@@ III.1 Band structure
Figure 4: Nonreciprocal heat current calculated by using Eq. (11) when (a) $\bm{h}=(0.1,-0.1,0.1)$, (b) $\bm{h}=(-0.1,0.1,0.1)$. At $T/J<0.10$, strong nonreciprocal current occurs. The change of magnetic direction inverts the direction of the current, which is corresponded to the band inversion.
We, first, show the band structures obtained by diagonalizing Eq. (10) in Figs. 3(a)-(d). In Figs. 3(a) and 3(b), the band structures are asymmetric, while, in Figs 3(c) and 3(d), the band structures are distorted by DMI but still symmetric. The degree of the asymmetry depends on the edge-parallel hoppings ($D_{1}$, $D_{3}$, $D_{5}$, $D_{6}$, $D_{7}$ and $D_{8}$), while the other terms ($D_{2}$ and $D_{4}$), i.e., $h_{z}$ do not affect it. Additionally, the band structures shown in Figs. 3(a) and 3(b) are related to each other via the inversion operation ($k_{X}\rightarrow-k_{X}$) because the sign changes of magnetic fields, “$h_{x}\rightarrow-h_{x}$ and $h_{y}\rightarrow-h_{y}$”, results in the change in the signs of the edge-parallel hoppings. The band structures shown in Figs. 3(c) and 3(d) are the same, when the hoppings $D_{1}$ and $D_{3}$ are zero under the condition $h_{x}=h_{y}$. Notably, it is advantageous for $h_{x}$ and $h_{y}$ to have opposite signs to maximize edge-parallel hoppings and manifest the effects of the edges.
Figure 5: (a) Results of the calculation of $\kappa^{(1)}$ in Eq. (12). (b) Results of the calculation of $\kappa^{(2)}$ in Eq. (12). Each of blue, red and green line represents $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$. The sign is inverted between $\bm{h}=(-0.1,0.1,0.1)$ and $\bm{h}=(0.1,-0.1,0.1)$. In the low temperature ($T/J\lesssim 0.09$), $\kappa^{(2)}$ is larger than $\kappa^{(1)}$ in $\bm{h}=(-0.1,0.1,0.1)$ and $\bm{h}=(0.1,-0.1,0.1)$. The inset shows the same data at the extreme low temperature. The value oscillates with the respect to the temperature and reaches zero.
@@ III.2 Nonreciprocal current and conductivity
In Figs. 4(a) and 4(b), we show the color map of the heat current plotted as functions of temperature and temperature gradient. Note again that Eq. (11) includes the contributions from both upper and lower edges of the system, in contrast to previous studies, for example Ref. [8]. Thus, the nonreciprocal behaviors are not due to chiral character of the edge states, but arise from the asymmetric Majorana bands caused by the DMI. The magnetic fields $\bm{h}$ in Figs. 4(a) and 4(b) correspond to those in Figs. 3(a) and 3(b). The pronounced nonreciprocal heat transport is observed at low temperature ($T/J<0.10$), where the heat current flows in the one direction regardless of the heat gradient. It is important to note that other contributions to the heat current, such as phonon contributions, should be considered to quantitatively evaluate the nonreciprocity. However, we anticipate that the nonreciprocal heat transport mediated by Majorana edge modes can be clearly observed in the low temperature regime, despite considering contributions from both upper and lower edges. Indeed, even though $\kappa^{(2)}$ retains a finite value above $T\gtrsim 0.10$, the reciprocal part $\kappa^{(1)}$ increases to offset the nonreciprocal contribution. The direction of the current differs between Figs. 3(a) and 3(b) due to the sign flip of magnetic fields.
We next explore the potential to manipulate the direction and magnitude of nonreciprocal heat current by varying external magnetic fields. For a detailed analysis, the results of the heat conductivities, $\kappa^{(1)}$ and $\kappa^{(2)}$ given by Eq. (12), are shown in Figs. 5(a) and 5(b). Each blue, red and green line corresponds to $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$, respectively. In Fig. 5(a), blue and red lines of $\kappa^{(1)}$ are numerically identical. However, for $\kappa^{(2)}$ in Fig. 5(b), they have the same absolute value but the opposite sign due to the combined influence of the band distortion and orientation, governed by $v^{3}$ in Eq. (12). The majority of contributions to the nonlinear conductivity come from the Majorana edge bands, as inversion symmetry is broken only at both edges in our model. In fact, $\kappa^{(2)}$ for $\bm{h}=(0.1,0.1,0.1)$ (shown as a green line in Fig. 5(b)) is numerically zero because the band structure is symmetric.
@@
Figure 6: (a) The direction of magnetic field dependance of $\kappa^{(2)}$. The nonlinear conductivity is maximized at $\theta=3\pi/4,7\pi/4$ and vanishes at $\theta=\pi/4,5\pi/4$. Each of blue, red and green line represents $T/J=0.06$, $T/J=0.1$ and $T/J=0.14$. (b) What the horizontal axis $\theta$ in (a) represents. We vary the magnetic field to keep $|\bm{h}|$ uniform.

### v2

@@ Abstract
Nonreciprocal transport, characterized by its direction-selective nature, holds significant potential for applications in various devices. In this study, we investigate nonreciprocal heat transport in Majorana systems, specifically focusing on the Kitaev chiral spin liquid under external magnetic fields. Our theoretical examination focuses on effects of open boundaries in which the Majorana edge modes exist, and the inversion symmetry is broken, which leads to the Dzyaloshinskii-Moriya interaction (DMI). Through perturbation theory, we demonstrate that DMI induces asymmetric hopping, resulting in the asymmetry of the Majorana band. The results of nonreciprocal heat currents are presented for various directions of external magnetic fields, and we discuss the relation between the current and the field-directions. The potential exists to manipulate both of the directions and magnitude of the nonreciprocal current by varying external magnetic fields and apply to heat transfer devices.
@@ I Introduction
where the conductivity $\chi$ and the external drive field $\mathcal{D}$ satisfy $\chi_{\xi}=\chi_{-\xi}$ and $\mathcal{D}_{\xi}=-\mathcal{D}_{-\xi}$. In particular, if the even-order conductivity is finite, the current magnitudes $|\mathcal{J}|$ differ between positive and negative directions. Nonreciprocal transport phenomena result from inversion and time-reversal symmetry breaking, which can cause higher-order effects that break Onsager reciprocal relations [1, 2, 3]. Thus, the difference in the group velocities at $\bm{k}$ and $\bm{-k}$, a typical higher-order effect, can induce nonreciprocal transport phenomena. Nonreciprocal transport phenomena have been extensively studied in diverse materials, such as superconductors [4, 5, 6, 7], antiferromagnets [8, 9], and topological insulators [10, 11]. However, its exploration in Majorana systems remains relatively limited [12].
@@
Several transport phenomena have been explored in the Kitaev chiral spin liquid. The itinerant Majorana fermions in the bulk act as carriers of spin [21, 22, 23, 24] and heat [25, 26]. Recently, the nonreciprocal heat transport in the Kitaev spin liquid by applying the staggard magnetic fields is discussed [12]. Additionally, it has been suggested that the itinerant Majorana fermions at the edge play an important role in the spin current [27]. However, the nonreciprocal heat transport for the case with the open edge is still unclear, where the non-dissipative heat current exist and the inversion symmetry is broken. Also, the Dzyaloshinskii-Moriya interaction (DMI) [28, 29], which is induced by the broken inversion symmetry, has a significant impact in the Kitaev honeycomb model [30, 31, 32, 33].
In this paper, we investigate the nonreciprocal heat current, which arises from chiral Majorana edge states, and can be manipulated by varying external magnetic fields. We conduct the calculation for the system which has the thermal gradient with consideration of the open edge [see Fig. 1(a)]. The unit cell is extended for the upper and lower sides to be the open boundaries, shown in Fig. 1(b). We consider the DMI at edges due to the broken inversion symmetry at open boundaries. Thereby, asymmetric hoppings and asymmetric Majorana bands appear. The heat current is evaluated by using Boltzmann transport theory. As a result, we find that the nonreciprocal heat current depends on the direction of the magnetic fields.
@@ II.1 Kitaev spin liquid with edge
where $S^{\gamma}_{j}$ is the $\gamma~{}(=x,y,z)$ component of an $S=1/2$ spin operator at site $j$. To avoid confusion, we adopt $x$, $y$ and $z$ as the orientations of the spin space, while $X$ and $Y$ denote the orientations of the real space, as shown in Fig. 1. The initial term in Eq. (2) represents the Kitaev interaction, characterized as an Ising-type interaction only for the nearest-neighbor (NN) sites. The notation $\langle jk\rangle_{\gamma}$ indicates the summation over NN sites connected by $\gamma$-bonds. Here, we set the interaction strength $J$ as the unit of energy with $J>0$, but our results are independent of the sign of $J$. The second term in Eq. (2) is a perturbation term induced by the Zeeman coupling $H_{\mathrm{h}}=-\sum_{j}h_{\gamma}S^{\gamma}_{j}$. The mean flux excitation energy is $\Delta$, and we set $\Delta=0.065J$. The coefficient $3!$ is due to the permutation of three spin operators arising from the perturbative processes. The notation $\langle\!\langle jkl\rangle\!\rangle$ denotes the summation exclusively over triplets of sites $j,k,$ and $l$ that are aligned and interconnected. Eq. (2) can be exactly solved using the well-known Majorana representation, $S^{\gamma}_{j}=\frac{i}{2}b^{\gamma}_{j}c_{j}$. Here, both $b_{j}^{\gamma}$ and $c_{j}$ satisfy the Majorana condition $\eta^{\dagger}=\eta$ and $\{\eta_{j}^{\alpha},\eta_{k}^{\beta}\}=\delta_{jk}\delta^{\alpha\beta}$ (with $\eta_{j}^{\alpha}=b_{j}^{x},b_{j}^{y},b_{j}^{z},c$). We start by considering a thermal current in the ground state, i.e., $u_{j,k}=ib^{\gamma}_{j}b^{\gamma}_{k}=1$ ($j\in{\textrm{even sublattice}}$, $k\in{\textrm{odd sublattice}}$) for all bonds. Each of the first and second terms in Eq. (2) transforms into a NN hopping and a next nearest-neighbor (NNN) hopping, respectively. When $h_{x}h_{y}h_{z}\neq 0$, the Chern number takes values $\nu=\pm 1$, reflecting the sign of $h_{x}h_{y}h_{z}$, and the non-dissipative chiral current is realized at the edge.
@@
Moreover, the DMI at the edges is taken into account due to the broken inversion symmetry at these boundaries. The DMI is expressed as,
@@
where the summation is restricted to NN sites ($j\in{\textrm{even sublattice}}$, $k\in{\textrm{odd sublattice}}$), with at least one site belonging to the edge. For simplicity, we assume that the inversion symmetry in the bulk is not broken, considering the DMI only at the edges. In the second-order perturbation combining the DMI and magnetic fields, the effective Majorana Hamiltonian is given as follows:
@@ II.3 Boltzmann transport equation
$v_{k_{X}n}$ represents the group velocity in the $X$ direction ($v_{k_{X}n}=\partial E_{k_{X}n}/{\partial k_{X}}$) and $n$ is the index of Majorana bands. The contributions from both the Majorana edge bands and the bulk are summed up in the second summation of Eq. (12). $\Omega$ denotes the system size, $\Omega=Nk_{\mathrm{mesh}}$, where $k_{\mathrm{mesh}}=1/N_{\mathrm{unit}}$. (A unit of length is a lattice constant.) It is important to note that the energy current is equivalent to the heat current, since the chemical potential of Majorana particles can be set to zero [25]. We adopt the Boltzmann constant value $k_{\mathrm{B}}=1$, and $f$ represents Fermi distribution function, $f(E_{k_{X}n})=(\exp(E_{k_{X}n}/T)+1)^{-1}$. Here, $\tau$ is the relaxation time. Assuming that backscattering of Majorana edge modes occurs mediated by the bulk excitations at finite temperature, we employ the relaxation time approximation and set $J\tau=1$. The second-order conductivity is nonlinear conductivity, which provides the nonreciprocal contribution.
@@ III.1 Band structure
Figure 4: (a) The nonreciprocity of the current, which is represented by the difference in heat currents for the oppositely applied thermal gradients, $|\mathcal{J}(\partial T/\partial X=1)|-|\mathcal{J}(\partial T/\partial X=-1)|$. The blue, red and green lines correspond to $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$, respectively. Changing magnetic direction inverts the direction of the current, corresponding to the band inversion. (b) The color map of the nonreciprocal heat current focused on the low temperature range for $\bm{h}=(-0.1,0.1,0.1)$. At $T/J<0.10$, strong nonreciprocal current occurs.
We, first, show the band structures obtained by diagonalizing Eq. (10) in Figs. 3(a)-(d). In Figs. 3(a) and 3(b), the band structures are asymmetric, while, in Figs 3(c) and 3(d), the band structures are distorted by the DMI but still symmetric. The degree of the asymmetry depends on the edge-parallel hoppings ($D_{1}$, $D_{3}$, $D_{5}$, $D_{6}$, $D_{7}$ and $D_{8}$), while the other terms ($D_{2}$ and $D_{4}$), i.e., $h_{z}$ do not affect it. Additionally, the band structures shown in Figs. 3(a) and 3(b) are related to each other via the inversion operation ($k_{X}\rightarrow-k_{X}$) because the sign changes of magnetic fields, “$h_{x}\rightarrow-h_{x}$ and $h_{y}\rightarrow-h_{y}$”, results in the change in the signs of the edge-parallel hoppings. The band structures shown in Figs. 3(c) and 3(d) are the same, when the hoppings $D_{1}$ and $D_{3}$ are zero under the condition $h_{x}=h_{y}$. Notably, it is advantageous for $h_{x}$ and $h_{y}$ to have opposite signs to maximize edge-parallel hoppings and manifest the effects of the edges.
Figure 5: (a) Results of the calculation of $\kappa^{(1)}$ in Eq. (12). The red broken line overlaps with the blue line. (b) Results of the calculation of $\kappa^{(2)}$ in Eq. (12). Each of blue line, red broken line and green line represents $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$. For $\kappa^{(2)}$, the sign is inverted between $\bm{h}=(-0.1,0.1,0.1)$ and $\bm{h}=(0.1,-0.1,0.1)$. In the low temperature ($T/J\lesssim 0.09$), $\kappa^{(2)}$ is larger than $\kappa^{(1)}$ in $\bm{h}=(-0.1,0.1,0.1)$ and $\bm{h}=(0.1,-0.1,0.1)$. The inset shows the same data at the extreme low temperature. The value oscillates with the respect to the temperature and reaches zero.
@@ III.2 Nonreciprocal current and conductivity
In Fig. 4(a), we show the nonreciprocity of the current, defined as the difference in heat currents for the oppositely applied thermal gradients, $|\mathcal{J}(\partial T/\partial X=1)|-|\mathcal{J}(\partial T/\partial X=-1)|$. Each of blue, red and green line represents $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$. The blue and red lines exhibit the same magnitude with opposite signs, reflecting the asymmetry and the inversion of the band structures shown in Figs. 3(a) and 3(b). Additionally, the nonreciprocity has two peaks at $T/J\simeq 0.1$ and $0.2$. In contrast, the green line shows no nonzero value, corresponding to the symmetric band structure in Fig. 3(c). In Fig. 4(b), we show the color map of the heat current plotted as functions of temperature and temperature gradient. Note again that Eq. (11) includes the contributions from both upper and lower edges of the system, in contrast to previous studies, for example Ref. [11]. Thus, the nonreciprocal behaviors are not due to chiral character of the edge states, but arise from the asymmetric Majorana bands caused by the DMI. The magnetic fields $\bm{h}$ in Fig. 4(b) correspond to those in Fig. 3(b) and the blue line in Fig. 4(a). The nonreciprocity reaches its peak at $T/J\simeq 0.1$ in Fig. 4(a), where the value of $|\mathcal{J}(\partial T/\partial X=1)|$ turns negative. The pronounced nonreciprocal heat transport is observed at low temperature ($T/J<0.10$), where the heat current flows in the one direction regardless of the heat gradient. It is important to note that other contributions to the heat current, such as phonon contributions, should be considered to quantitatively evaluate the nonreciprocity. However, we anticipate that the nonreciprocal heat transport mediated by Majorana edge modes can be clearly observed in the low temperature regime, despite considering contributions from both upper and lower edges. Indeed, even though $\kappa^{(2)}$ retains a finite value above $T\gtrsim 0.10$, the reciprocal part $\kappa^{(1)}$ increases to offset the nonreciprocal contribution. The direction of the current differs between $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(-0.1,0.1,0.1)$ (same as Fig. 4(b)), corresponding to the red and blue line in Fig. 4(a).
@@
Figure 6: (a) The direction of magnetic field dependence of $\kappa^{(2)}$. The nonlinear conductivity is maximized at $\theta=3\pi/4,7\pi/4$ and vanishes at $\theta=\pi/4,5\pi/4$. Each of blue, red and green line represents $T/J=0.06$, $T/J=0.1$ and $T/J=0.14$. (b) What the horizontal axis $\theta$ in (a) represents. We vary the magnetic field to keep $|\bm{h}|$ uniform.
We next explore the potential to manipulate the direction and magnitude of nonreciprocal heat current by varying external magnetic fields. For a detailed analysis, the results of the heat conductivities, $\kappa^{(1)}$ and $\kappa^{(2)}$ given by Eq. (12), are shown in Figs. 5(a) and 5(b). Each blue line, red broken line and green line corresponds to $\bm{h}=(-0.1,0.1,0.1)$, $\bm{h}=(0.1,-0.1,0.1)$ and $\bm{h}=(0.1,0.1,0.1)$, respectively. In Fig. 5(a), blue line and red broken line of $\kappa^{(1)}$ are numerically identical. However, for $\kappa^{(2)}$ in Fig. 5(b), they have the same absolute value but the opposite sign due to the combined influence of the band distortion and orientation, governed by $v^{3}$ in Eq. (12). The majority of contributions to the nonlinear conductivity come from the Majorana edge bands, as inversion symmetry is broken only at both edges in our model. In fact, $\kappa^{(2)}$ for $\bm{h}=(0.1,0.1,0.1)$ (shown as a green line in Fig. 5(b)) is numerically zero because the band structure is symmetric.